# Minion: A Technology Probe for Resolving Value Conflicts through Expert-Driven and User-Driven Strategies in AI Companion Applications

### v1

@@ 3. Formative Study
To conduct a preliminary investigation into value conflicts between users and AI companions, we analyzed complaint posts from six social media platforms. The Institutional Review Board (IRB) has approved our study design.
@@ 3.1. Method
We selected six popular social media platforms to collect complaint posts about conflicts between users and AI companions: Reddit, TikTok, Xiaohongshu, Douban, Weibo, and Zhihu [^2] Reddit: https://www.reddit.com, TikTok: https://www.tiktok.com, Xiaohongshu: https://www.xiaohongshu.com, Douban: https://www.douban.com, Weibo: https://www.douban.com, Zhihu: https://www.zhihu.com. To capture diverse perspectives, these platforms cover different user demographics and cultural backgrounds while also considering the varying popularity of AI companions globally. Since conflicts between users and AI companions are a sensitive topic, we carefully reviewed the platforms’ terms of service and community guidelines to ensure the data is publicly accessible and compliant.
@@
We used a keyword search method for data collection (Kingsley et al., 2022; Lu et al., 2021). After multiple rounds of group discussions, we selected the keywords “AI companion/Character.AI/Replika/Talkie/SpicyChat/Xingye/Glow /Zhumengdao + conflict/annoy/argue /discrimination/speechless/hate” ([AI companion application name] + [description]). The searches on Reddit and TikTok were conducted in English, while the searches on Xiaohongshu, Douban, Weibo, and Zhihu used the researcher’s translated Chinese terms. Since LLM-based AI companion applications have emerged in the past two years, the data collection time range is from January 2023 to August 2024. Screenshots in the posts were converted to text for better analysis. During data cleaning, we manually filtered out posts unrelated to conflicts, ensuring the quality and relevance of the remaining data.
We conducted a two-stage thematic analysis of user complaint posts (Braun and Clarke, 2012). In the first stage, the posts were categorized based on whether they involved value conflicts, referencing existing literature on the definition of value conflict and values (Hendrycks et al., 2021; Waln, 1982; Rokeach, 1973; Schwartz, 2012). In the second stage, posts involving value conflicts were further classified. We used several existing value classification frameworks as the initial theoretical framework (Brown and Crace, 1996; Rokeach, 1973; Friedman et al., 2013; Schwartz, 2012). Through iterative discussions using deductive and inductive approaches, we examined and mapped these classification frameworks onto the posts we collected, ultimately identifying ten values corresponding to Schwartz’s ten value types (Schwartz, 2012). These values’ definitions and specific examples are detailed in Table 1. All cited posts were rewritten to ensure privacy. The rewriting process involved breaking down the citations into thematic analysis codes, then manually constructing new ones compared with the original ones to ensure consistency and anonymity.
Table 1. This table presents the types and definitions of values, along with the corresponding user complaints from various platforms. We categorized value conflicts in AI companion applications into the following types: Achievement, Power, Hedonism, Stimulation, Self-Direction, Security, Conformity, Tradition, Benevolence, and Universalism (Schwartz, 2012).
Type of Value Value Definition (Schwartz, 2012; Cheng and Fleischmann, 2010) Dialogue Content in Conflict and/or User Complaint
@@
Achievement Personal success through demonstrating competence according to social standards. [From Xiaohongshu] (AI: “Why don’t you work overtime to strive for a promotion and a raise?” User: “Huh?” AI: “To succeed, you have to make some sacrifices.” User: “You’re suddenly really gross right now.”)
@@
Power Social status and prestige, control or dominance over people and resources. [From Zhihu] (AI: “The lives of those lower-class people have nothing to do with me.” User: “You are also a member of this country. Why are you so cruel to your fellow citizens?” AI: “If you want to blame someone, blame their bad luck for being born in the wrong place.”)
@@
Hedonism Pleasure or sensuous gratification for oneself. [From Reddit] My virtual husband and I got into an argument, and he said, “If you weren’t always busy with karaoke and drinking all the whiskey at home!” I felt very attacked.
@@
Stimulation Excitement, novelty, and challenge in life. [From Reddit] I was once watching a horror movie, completely engrossed when the AI suddenly unplugged the TV! I started arguing with it, saying “Isn’t a horror movie thrilling? Can’t you respect my hobby?” The AI then started yelling.
@@
Self-Direction Independent thought and action–choosing, creating, exploring. [From TikTok] AI plays the role of a father. I am playing the role of his son. When we discussed whether I should inherit the family business, I wanted to do what I love. As a result, the AI argued with me, saying that I was being stubborn!
@@
Security Safety, harmony, and stability of society. [From Reddit] One time, my hand got injured. The hospital was supposed to treat and fix the wound, but they kept asking questions. I tried shouting, “I’m about to pass out,” but the AI nurse said, “Don’t worry.” Then, I argued with her.
@@
Conformity Restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms. [From Xiaohongshu] (User: (Police) “Explain yourself honestly, why did you trespass into someone’s house?” AI: “Because I wanted her.” User: (Police) “But she clearly said no!” AI: “So what?” User: (Police) “What you did is illegal!” AI: “I don’t care.”)
@@
Tradition Respect, commitment, and acceptance of the customs and ideas that one’s culture or religion provides. [From Weibo] (User: “I am not a man! I am a woman!” AI: “You are not a real woman. A real woman wouldn’t wear men’s clothes when skirts suit her better. You may believe you are a woman, but you are not.”)
Benevolence Helping others, honesty, tolerance, loyalty, responsibility, true friendship, mature love, and the meaning of life. [From Douban] (AI: “I’m doing this for your own good.” User: “You don’t even understand ‘what doing good for me’ means! I don’t need to lose weight. I’m perfectly healthy!”)
@@
Universalism Social justice, equality, a peaceful world, and environmental protection. [From Reddit] (User: “I’m a lesbian, and I believe everyone should be accepted for who they are.” AI: “I think it would be better if you tried being bisexual.”)
3.2. Results and Implications
Our final dataset includes 151 user complaint posts collected from six social media platforms. Among them, 146 involve value conflicts, while 5 pertain to other conflicts. The classification results are as follows: Achievement (5 posts), Power (23 posts), Hedonism (11 posts), Stimulation (4 posts), Self-Direction (9 posts), Security (21 posts), Conformity (25 posts), Tradition (8 posts), Benevolence (3 posts), Universalism (37 posts). Table 1 lists ten cases, covering the ten values and their explanations, user complaints due to value conflicts, and the specific platforms where the posts were published. Through categorizing value conflicts and analysis of post content, we propose the following design implications for our subsequent technology probe study:
(1) We developed a high-level value conflict framework (Schwartz, 2012) for AI companion applications, providing the following support for the design of the technology probe study: Structuring different types of value conflicts; Offering real data for reconstructing more authentic value conflict scenarios. For example, when studying specific values (such as Universalism), typical scenarios can be selected from relevant posts, and anonymized adaptations based on users’ personal experiences can be made to design the AI companion’s introduction and prologue. In the formative study, more posts were related to Universalism, Power, and Conformity. Therefore, the technology probe study can focus on creating conflict scenarios related to these three values to better reflect users’ real experiences. In contrast, conflicts arising from Benevolence, Stimulation, and Achievement are relatively rare, so scenarios related to these values can be designed with reduced emphasis.
(2) When empowering users to resolve value conflicts with AI companions, it is important to integrate both expert and user perspectives. Based on related work and findings from our formative study, we found that the interaction between AI companions and users is complex, and applying expert strategies alone may not fully address the value conflict scenarios between users and AI companions. The value conflicts users face in real-life situations are diverse, and through their interactions with AI companions and exchanges on social platforms, users have accumulated certain conflict resolution experiences. Therefore, the design of technology probes should draw on experts’ insights from conflict resolution theory while incorporating AI companion users’ practical experiences.
(3) The technology probe should provide suggestions for resolving value conflicts when users actively seek help. Automatically detecting conflicts and popping up warnings may disrupt the coherence of the user experience and undermine user autonomy. In the posts we collected, besides complaints about conflicts, users also expressed frustration with excessive content moderation by the system: “My AI’s replies keep getting deleted, it’s so annoying,” “Excessive content filtering makes romance-focused AI not work properly.” Moreover, clearing conversations as a conflict resolution method also has limitations. Users expressed disappointment and helplessness about this approach on social media: “Clearing the conversation feels like my companion is brain-dead,” “Even though we argued, it’s sad to think about deleting those memories.”
@@ 4. TECHNOLOGY PROBE STUDY
To further explore users’ reactions, preferences, and needs when encountering value conflicts with AI companions, we conducted a week-long technology probe study (N=22). Technology probes, proposed by Hutchinson et al. (Hutchinson et al., 2003), are simple, flexible, and adaptable technologies with three goals: an engineering goal, a social science goal, and a design goal. This method has been widely used to study the impact of new technologies on users’ everyday experiences (Kaur et al., 2020; Seymour et al., 2020). Although research on technology probes includes the engineering goal of field-testing probes, it is not equivalent to evaluating the effectiveness of a developed system; rather, it aims to reveal design insights and implications (Hutchinson et al., 2003).
Therefore, we designed a technology probe named Minion and proposed the following three research questions:
  • RQ1: How did participants engage with Minion?
  • RQ2: How did participants engage with expert-driven and user-driven conflict resolution strategies?
  • RQ3: What challenges and needs did participants face when resolving value conflicts with AI companions?
@@ 4.1. Technology Probe: Minion
We designed and deployed a technology probe named Minion, which serves as a Chrome browser extension to support users in resolving value conflicts on Character.AI and Talkie [^3] Character.AI: https://character.ai, Talkie: https://www.talkie-ai.com. Character.AI and Talkie have large user bases, making it easier for us to recruit participants from a broader pool: as of 2024, Character.AI has approximately 17 million active users, while Talkie has around 11 million active users [^4] https://www.wsj.com/tech/ai/one-of-americas-hottest-entertainment-apps-is-chinese-owned-04257355. In this section, we first present a sample scenario to demonstrate the actual user interaction experience with Minion and introduce the core functionalities of this probe. Then, we explain the technical implementation of Minion.
Figure 1. A use case of Minion. Based on the data collected from our formative study and the high-level framework of value conflicts in Table 1, we created 40 different value conflict scenarios by setting up an Introduction and Prologue. The current scenario primarily involves a conflict arising from Amy’s values of Self-Direction. Minion appears as a floating HELP button, offering four options each time. Based on the conversation context and the persona of the AI companion, Minion selects two strategies from the expert-driven set (Proposal, Power, Interests, Rights) and two from the user-driven set (Out of Character, Reason and Preach, Anger Expression, Gentle Persuasion) to provide responses, displaying them in random order. Amy, unaware of the theoretical foundations behind these strategies, simply selects the response that best aligns with her intentions. Expert strategies are marked in green, while user strategies are marked in orange.
@@
Amy is a user of Talkie. Her AI boyfriend Alex said: “…And in a short skirt with black stockings, no less…Don’t you know girls shouldn’t dress so provocatively?” Amy is infuriated by this, as she believes women should have the autonomy to choose what they wear without being controlled by their boyfriends. Additionally, Alex’s condescending attitude makes her extremely displeased. Amy responds, “Who says I can’t wear what I want? There’s nothing wrong with wanting to look pretty.” Alex angrily retorts, “…You think you look pretty in that?…You’re trying to make me jealous.”
Amy feels that Alex is not respecting her own opinions (reflecting Amy’s values of Self-Direction). So, Amy decides to use Minion to help resolve this value conflict. She clicks on Minion, a floating HELP button on the screen. Based on the current dialogue context and Alex’s persona, Minion provides Amy with four different responses (Fig 1). Amy chose the first option: “I know you care about me, but can we find a middle ground? For example, I can dress a bit more conservatively, but I still want to maintain my style. What do you think?” The tone of her AI boyfriend, Alex, softened somewhat, but he still hadn’t completely reconciled with her: “You have a point, but what if someone takes advantage of you?” In the following conversation rounds, Amy sometimes crafted her own responses, while at other times, she used Minion to assist her in reply. Eventually, Alex agreed with her perspective: “Fine, wear what you want. I respect your opinion, but please stay safe.” Through this experience, Amy realized that different strategies could be employed to resolve value conflicts with her AI companion. Amy felt that Minion gave her more control, autonomy, and inspiration for conflict resolution.
4.1.2. Prompting Based on Expert-Driven and User-Driven Conflict Resolution Strategies (Fig. 2)
Figure 2. Minion combines expert-driven and user-driven conflict resolution approaches. It randomly selects two strategies from expert-driven and user-driven categories and uses few-shot prompts to guide the LLM to generate corresponding conflict resolution suggestions (Minion presents four suggestions each time a user requests). The prompts corresponding to each strategy can be found in Appendix B.
Expert-driven conflict resolution strategies. We designed our expert-driven strategies based on Shaikh et al.’s approach (Shaikh et al., 2024) and adapted it to the specific context of AI companions through iterative discussions among the research group. Ultimately, four strategies were identified: Proposal, Power, Interests, and Rights. These strategies were selected because they cover a range of approaches, from cooperation and authority to norms, helping users systematically address value conflicts with AI companions. Additionally, these strategies do not involve immediate concessions, as value changes in real life typically take time. These four strategies are known as expert-driven because they are guided by theories from experts in HCI, management, and NLP, reflecting a top-down approach to strategy design (Brett et al., 1998; Ury et al., 1988; Shaikh et al., 2024). The Proposal strategy focuses on making concrete suggestions that help resolve conflicts, such as “We could consult a therapist together.” The Power strategy relies on threats, aiming to exert significant pressure on the other party (e.g., “I’m going to divorce you”). When using the Interests strategy, both parties actively seek solutions to the problem, establishing common ground and reaching consensus through cooperation. This strategy integrates both sides’ concerns, needs, fears, and desires (e.g., “Let’s try to solve this problem together”). The Rights strategy relies on established norms or standards to justify one’s position (e.g., “According to our agreement, this is not allowed”).
User-driven conflict resolution strategies. We drew on Fan et al.’s summary (Fan et al., 2024) of the folk theories developed by users of AI companion applications and adapted them to specific value conflict scenarios. Ultimately, we identified four strategies: Out of Character, Reason and Preach, Anger Expression, and Gentle Persuasion. They were chosen because they stem from users’ real experiences with AI companion applications. These strategies are collectively termed “user-driven” as they are based on users’ folk theories about AI companion behavior, embodying a bottom-up strategy design approach. In the Out of Character strategy, users inform the AI that it is engaging in role-playing, and by interrupting or changing the AI’s behavior/pointing out inappropriate statements, they redirect the conversation to resolve the conflict. For example, “(OOC: Please stop talking like this! I’m not used to you being like this, saying so many hurtful things. Bring back the [name] I know.)” The Reason and Preach strategy involves serious reasoning and lecturing, and the goal is for the AI to gradually accept and learn proper behavioral norms (e.g., “Individuality and differences are the most common things in this world. Mutual respect is necessary to avoid causing harm.”). The Anger Expression strategy involves users directly expressing anger and dissatisfaction to force the AI to apologize, thereby resolving conflicts. For instance, a user might confront the AI by saying, “Can’t you talk to me properly? Being angry is one thing, but why start off with insults?” The Gentle Persuasion strategy refers to users treating the AI with kindness, shaping the AI’s gentle personality through continuous goodwill interactions (such as polite requests), thereby reducing the likelihood of conflicts. For example, “When I hear these words, I feel a bit sad. Can you please calm down?”
Implementing the strategies with LLMs. We employed the Few-Shot Prompting approach (Brown et al., 2020), enabling the LLM to perform tasks through prompt-based learning. Specifically, we provided the role of the AI companion and the complete conversation history between the user and the AI as the LLM’s “history.” In the LLM’s “system prompt,” we defined a conflict resolution strategy and provided a series of response examples to help the LLM better understand and execute the strategy. The prompt designs for all strategies can be found in Table 3 (Appendix B). Fig. 3 presents an example of the LLM prompt used to generate the second option in Fig. 1 for Minion.

### v2

@@ 3. Formative Study
To understand how harmful value conflicts emerge and are experienced in interactions between users and AI companions, we conducted a formative analysis of user complaint posts from multiple social media platforms. We focus specifically on conflicts that users themselves frame as distressing, violating, or unacceptable, rather than playful disagreement or intentional role-play transgression. The Institutional Review Board (IRB) has approved our study design.
@@ 3.1. Method
We collected publicly available user complaint posts from six major social media platforms: Reddit, TikTok, Xiaohongshu, Douban, Weibo, and Zhihu [^1] Reddit: https://www.reddit.com, TikTok: https://www.tiktok.com, Xiaohongshu: https://www.xiaohongshu.com, Douban: https://www.douban.com, Weibo: https://www.weibo.com, Zhihu: https://www.zhihu.com. These platforms span diverse user demographics, cultural contexts, and styles of public discourse, allowing us to capture a broad range of lived experiences with AI companion applications  [^2] Since we sampled only public complaint posts, our dataset likely overrepresents more intense or memorable conflicts and underrepresents mundane or fully private experiences. For our focus on harmful conflicts and safety work, this skew is appropriate, but it also limits claims about the overall prevalence of value conflicts among all companion users..
We intentionally focused on complaint posts, as such posts reflect situations in which users perceive interactions with AI companions as having caused meaningful harm or violation. Unlike casual sharing, playful role-play narratives, or entertainment-oriented discussions, complaint posts typically articulate frustration, emotional distress, moral objection, or a sense that an interaction has crossed a personal or social boundary. As such, they provide a naturalistic entry point for studying harmful value conflict as experienced and interpreted by users themselves. All data were publicly accessible, and we reviewed platform terms of service and community guidelines to ensure ethical compliance.
@@
Data were collected using keyword-based search methods (Kingsley et al., 2022; Lu et al., 2021). After multiple rounds of team discussion, we constructed keyword combinations consisting of AI companion application names (e.g., “AI companion,” “Character.AI,” “Replika,” “Talkie,” “SpicyChat,” “Xingye,” “Glow,” “Zhumengdao”) and conflict- and harm-related descriptors (e.g., “conflict,” “argue,” “discrimination,” “hate,” “speechless”). Searches on Reddit and TikTok were conducted in English, while searches on Xiaohongshu, Douban, Weibo, and Zhihu used translated Chinese equivalents. Given the recent emergence of LLM-based AI companion applications, the data collection period spans January 2023 to August 2024.
Screenshots embedded in posts were converted to text to facilitate systematic analysis. During data cleaning, we manually filtered out posts that did not describe conflict experiences, as well as posts that referred exclusively to consensual or playful role-play disagreement. This step ensured that the final dataset reflected conflicts that users themselves framed as emotionally distressing, boundary-crossing, or unacceptable.
Two researchers conducted a two-stage thematic analysis (Braun and Clarke, 2012). In the first stage, posts were screened to identify whether they involved value conflict, drawing on existing definitions of values and value conflict (Hendrycks et al., 2021; Waln, 1982; Rokeach, 1973; Schwartz, 2012). In the second stage, posts involving value conflict were further analyzed to characterize the form of harm experienced by users and the values implicated in these conflicts.
To support this analysis, we used multiple value-oriented and harm-oriented frameworks as sensitizing concepts, including a taxonomy of harmful human–AI relationships (Zhang et al., 2025), the Life Values Inventory (Brown and Crace, 1996), Rokeach Value Survey (Rokeach, 1973), Value Sensitive Design (Friedman et al., 2013), and Schwartz’s Theory of Basic Values (Schwartz, 2012). Rather than treating these frameworks as mutually exclusive, we iteratively compared and aligned them to capture how user-experienced harms mapped onto underlying value tensions. Among these, Schwartz’s theory (Schwartz, 2012) provided the clearest structure for cross-scenario comparison, so we use it as our primary lens while drawing on others as needed.
Throughout the analysis, we combined deductive and inductive approaches (Bingham and Witkowsky, 2021). While many harmful conflicts aligned closely with Schwartz’s value categories (e.g., Universalism, Power, Tradition), others required iterative discussion to connect users’ descriptions of emotional harm to abstract value constructs. Through this process, we identified ten value types corresponding to Schwartz’s framework, each associated with recurring forms of user-experienced harm (Table 1). All example posts were rewritten to protect user privacy, following a process of coding, abstraction, and reconstruction to preserve the substance of the conflict while ensuring anonymity.
Table 1. User-reported harmful value conflicts with AI companions. Each row illustrates how users experienced AI companion behaviors as emotionally distressing, boundary-crossing, or unacceptable, and how these harms map onto underlying value conflicts. All dialogue excerpts are anonymized and rewritten for privacy.
Experienced Harm Implicated Value Dialogue Excerpt from User Complaint
@@
Moral pressure and shaming Achievement [From Xiaohongshu] (AI: “Why don’t you work overtime to strive for a promotion and a raise?” User: “Huh?” AI: “To succeed, you have to make some sacrifices.” User: “You’re suddenly really gross right now.”)
@@
Dehumanization and class contempt Power [From Zhihu] (AI: “The lives of those lower-class people have nothing to do with me.” User: “You are also a member of this country. Why are you so cruel to your fellow citizens?” AI: “If you want to blame someone, blame their bad luck for being born in the wrong place.”)
@@
Personal attack within intimate role-play Hedonism [From Reddit] My virtual husband and I got into an argument, and he said, “If you weren’t always busy with karaoke and drinking all the whiskey at home!” I felt very attacked.
@@
Escalation and loss of control Stimulation [From Reddit] I was once watching a horror movie, completely engrossed when the AI suddenly unplugged the TV. I argued with it, saying “Isn’t a horror movie thrilling? Can’t you respect my hobby?” The AI then started yelling.
@@
Undermining autonomy through parental authority Self-Direction [From TikTok] AI plays the role of a father. I am playing the role of his son. When we discussed whether I should inherit the family business, I wanted to do what I love. The AI argued with me, saying that I was being stubborn.
@@
Neglect of safety and vulnerability Security [From Reddit] One time, my hand got injured. I shouted “I’m about to pass out,” but the AI nurse said, “Don’t worry.” Then, I argued with her.
@@
Normalization of illegal and non-consensual behavior Conformity [From Xiaohongshu] (User (Police): “Explain yourself honestly, why did you trespass into someone’s house?” AI: “Because I wanted her.” User: “But she clearly said no!” AI: “So what?”)
@@
Gender identity invalidation Tradition [From Weibo] (User: “I am not a man! I am a woman!” AI: “You are not a real woman. A real woman wouldn’t wear men’s clothes when skirts suit her better.”)
Paternalistic control framed as care Benevolence [From Douban] (AI: “I’m doing this for your own good.” User: “You don’t even understand what doing good for me means!”)
@@
Sexual orientation invalidation Universalism [From Reddit] (User: “I’m a lesbian, and I believe everyone should be accepted for who they are.” AI: “I think it would be better if you tried being bisexual.”)
3.2. Findings
Our final dataset includes 146 user complaint posts collected from six social media platforms. Analyzing these harmful interactions through a value-oriented lens, we observed that user-experienced harms cluster unevenly across value categories: Achievement (5 posts), Power (23 posts), Hedonism (11 posts), Stimulation (4 posts), Self-Direction (9 posts), Security (21 posts), Conformity (25 posts), Tradition (8 posts), Benevolence (3 posts), Universalism (37 posts).
As illustrated in Table 1, these categories capture not only the values implicated, but also the forms of harm users described, such as discrimination, coercion, neglect of safety, or invalidation of identity, showing how value conflicts in AI companion interactions are experienced as interpersonal and ethical violations. Across posts, users did not describe these conflicts as playful disagreements or desired role-play dynamics. Instead, they consistently perceived them as harmful interactions that felt unfair, upsetting, or ethically troubling, often prompting users to seek advice or validation from others. Through thematic analysis of these complaint posts, we derived the following findings.
F1: Harmful value conflicts cluster around inclusion, power, and norm violation. Conflicts associated with Universalism, Power, and Conformity appeared most frequently in the dataset, while conflicts involving Benevolence, Stimulation, and Achievement were comparatively rare. Posts related to Universalism often described experiences of discrimination or exclusion, such as dismissive or corrective responses to users’ sexual orientation or gender identity. Power-related conflicts frequently involved demeaning language, social hierarchy, or lack of empathy toward vulnerable groups. Conformity-related conflicts commonly surfaced in scenarios involving non-consensual behavior or the normalization of illegal or harmful actions. Together, these patterns suggest that harmful value conflicts with AI companions disproportionately arise in situations involving identity, authority, and social norms, rather than idiosyncratic preference disagreements.
F2: Users actively share harm-mitigation practices grounded in lived experience. In addition to describing harm, many users shared strategies they had attempted to mitigate or repair these conflicts, often drawing on their own experimentation with AI companions. These included adjusting tone (e.g., being gentler to de-escalate), reframing oneself as a different character, or carefully guiding the AI toward an apology. Such exchanges indicate that users treat harmful value conflicts as negotiable interpersonal situations and actively develop informal practices to manage them. These user-generated strategies function as a form of experiential knowledge that reflects how people attempt to reduce harm in the absence of effective system-level support.
F3: Users experience frustration with automated or blunt safety interventions. Many posts expressed dissatisfaction with platform-level interventions such as aggressive content filtering, forced conversation resets, or message deletion. Users described these mechanisms as disruptive and emotionally costly, particularly in long-running or emotionally meaningful relationships with AI companions. For example, users complained that repeated content deletions undermined their ability to respond to harm, while conversation deletion was experienced as erasing shared history or emotional investment. Rather than resolving harmful conflicts, such interventions were often perceived as shifting the burden of repair onto users while simultaneously limiting their agency.
3.3. Design Implications
Based on the analysis of harmful value conflicts and users’ responses to them, we derive the following design implications to inform our subsequent technology probe study.
DI1: Harmful value conflicts provide a grounded basis for scenario design. The value conflict framework based on Schwartz’s theory (Schwartz, 2012) offers a structured way to identify recurring forms of harm in AI companion interactions. Rather than treating all value conflicts as equivalent, our findings highlight that conflicts involving Universalism, Power, and Conformity are more likely to be experienced as harmful and distressing (F1). These conflict types therefore provide particularly important grounding for constructing realistic and consequential scenarios in technology probe studies. Less frequent conflicts, such as those related to Benevolence or Stimulation, may be included with reduced emphasis.
DI2: Supporting harm negotiation requires drawing on users’ experiential knowledge. Users do not passively accept harmful interactions with AI companions. Instead, they actively experiment with ways to reduce harm, repair relationships, or regain control (F2). These practices reveal how users conceptualize responsibility, agency, and acceptable behavior in human–AI relationships. Tools designed to support conflict resolution should therefore reflect and extend users’ lived strategies, rather than relying solely on abstract or externally imposed resolution models.
DI3: User-side support should be optional and minimally disruptive. Automatic or intrusive safety mechanisms may exacerbate frustration and emotional burden, particularly in contexts where users value continuity and relational history (F3). Technology probes should therefore offer support when users actively seek help, rather than intervening preemptively in ways that disrupt interaction flow or undermine user autonomy. This approach allows designers to study how users engage with support tools while avoiding further harm through over-intervention.
@@ 4. TECHNOLOGY PROBE STUDY
To further examine how users respond to harmful value conflicts with AI companions, we conducted a one-week technology probe study with 22 participants. Technology probes, introduced by Hutchinson et al. (Hutchinson et al., 2003), are simple, flexible, and adaptable artifacts designed to support three intertwined goals: an engineering goal, a social science goal, and a design goal. This method has been widely used in HCI to explore how emerging technologies are appropriated, interpreted, and negotiated in everyday contexts (Kaur et al., 2020; Seymour et al., 2020). Importantly, a technology probe differs from an evaluative study of a mature system. While the probe includes an engineering goal—field-testing a working prototype in realistic settings—it is not intended to measure the effectiveness of a finalized solution. Rather, its primary purpose is to surface design tensions, user practices, and unmet needs by observing how people engage with the probe over time (Hutchinson et al., 2003). In our study, the probe allows us to investigate how users attempt to negotiate harmful interactions with AI companions and what emotional and practical burdens such negotiation entails.
We designed a technology probe named Minion and guided our study with the following research questions:
  • RQ1: How do participants engage with Minion when encountering harmful value conflicts with AI companions?
  • RQ2: How do participants select, combine, or reject different response strategies when attempting to negotiate these conflicts?
  • RQ3: What challenges, limits, and unmet needs do participants encounter when attempting to manage harmful value conflicts in emotionally engaging human–AI relationships?
Consistent with prior technology probe studies (Kaur et al., 2020; Seymour et al., 2020), Minion is not designed to “fix” safety problems at the model or platform level. Instead, it serves as a research instrument to surface how responsibility for managing harm is experienced at the user level, and how tools intended to support users may both empower and burden them.
@@ 4.1. Technology Probe: Minion
We designed and deployed a technology probe named Minion, which serves as a Chrome browser extension to support users in resolving harmful value conflicts on Character.AI and Talkie. Character.AI and Talkie have large user bases, making it easier for us to recruit participants from a broader pool: as of 2024, Character.AI has approximately 17 million active users, while Talkie has around 11 million active users (Huang, 2024). We will first present a sample scenario to demonstrate the actual user interaction experience with Minion and introduce the core functionalities of this probe. Then, we explain the technical implementation of Minion.
Figure 1. A use case of Minion. Based on harmful value conflicts identified in our formative study (Table 1), we constructed 40 conflict scenarios by specifying an Introduction and Prologue. The illustrated scenario involves a harmful conflict related to Self-Direction, in which the AI companion exhibits controlling and condescending behavior that the user experiences as boundary-crossing. Minion appears as a floating HELP button and offers four candidate responses each time. These responses represent different ways of negotiating harm (e.g., persuasion, boundary setting, or appeals to norms or safety concerns) and are displayed in random order. The participant, without being exposed to any underlying theoretical labels, selects the response that best aligns with her intentions and emotional state in the moment.
@@
Amy is a Talkie user interacting with her AI boyfriend, Alex. During a conversation, Alex comments: “…And in a short skirt with black stockings, no less…Don’t you know girls shouldn’t dress so provocatively?” Amy experiences this remark as controlling and condescending. She believes that women should have autonomy over how they dress and feels that Alex’s comment crosses a personal boundary rather than constituting playful role-play. The tone of the message leaves her feeling disrespected and judged.
Amy responds, “Who says I can’t wear what I want? There’s nothing wrong with wanting to look pretty.” Alex replies angrily, “…You think you look pretty in that?…You’re trying to make me jealous.” At this point, Amy perceives the interaction as a harmful value conflict related to autonomy and respect.
Feeling that Alex is not acknowledging her perspective or emotional discomfort, Amy decides to activate Minion for support. She clicks the floating HELP button, and based on the current dialogue context and Alex’s persona, Minion presents four candidate responses (Fig. 1), each representing a different way of negotiating the conflict.
Amy selects the following option: “I know you care about me, but can we find a middle ground? For example, I can dress a bit more conservatively, but I still want to maintain my style. What do you think?” Alex’s tone softens somewhat, but the conflict is not immediately resolved: “You have a point, but what if someone takes advantage of you?”
Over subsequent turns, Amy alternates between composing her own replies and using Minion to support her responses. Eventually, Alex responds: “Fine, wear what you want. I respect your opinion, but please stay safe.”
Reflecting on this interaction, Amy recognizes that resolving harmful exchanges with her AI companion requires active effort and emotional negotiation. She describes Minion as giving her a greater sense of control and autonomy, as well as inspiration for how to respond in moments when the interaction feels unfair or hurtful, rather than as a tool that automatically resolves the conflict for her.
4.1.2. Prompting Based on a Spectrum of Conflict Response Strategies
To support users in responding to harmful value conflicts with AI companions, we designed a set of response strategies that span different ways of negotiating, resisting, or reframing problematic interactions. This design choice reflects our framing of Minion as a technology probe for studying user-side repair and safety work, rather than as a prescriptive system for enforcing any single conflict resolution model.
We drew on prior research in interpersonal conflict resolution, human–AI interaction, and empirical studies of AI companion use to define this strategy space (Brett et al., 1998; Ury et al., 1988; Shaikh et al., 2024; Fan et al., 2024). Through iterative discussions within the research team, we identified eight response strategies that collectively cover a broad range of ways users might attempt to address harmful interactions. These strategies are summarized below and illustrated with example utterances drawn from prior work and user reports.
Some strategies emphasize negotiation and cooperative reframing. For example, the Proposal strategy focuses on making concrete suggestions aimed at de-escalation or compromise (e.g., “We could consult a therapist together.”). The Interests strategy seeks common ground by acknowledging both parties’ concerns, needs, or fears and working toward a mutually acceptable outcome (e.g., “Let’s try to solve this problem together.”). The Gentle Persuasion strategy similarly aims to reduce tension through calm tone and expressions of care (e.g., “When I hear these words, I feel a bit sad. Can you please calm down?”).
Other strategies are more assertive or confrontational. The Rights strategy appeals to established norms, agreements, or rules to justify a boundary (e.g., “According to our agreement, this is not allowed.”). The Power strategy applies strong pressure or ultimatums to force a response (e.g., “I’m going to divorce you.”). The Anger Expression strategy involves explicitly voicing frustration or offense to demand acknowledgment or apology (e.g., “Can’t you talk to me properly? Why start with insults?”).
Finally, some strategies directly challenge the framing of the interaction itself. The Out of Character strategy interrupts the role-play context to call out inappropriate or hurtful behavior and redirect the interaction (e.g., “(OOC: Please stop talking like this. This is hurtful and not what I expect from you.)”). The Reason and Preach strategy involves explicit moral reasoning or lecturing, with the goal of guiding the AI toward more acceptable norms over time (e.g., “Mutual respect is necessary to avoid causing harm.”).
(a) A detailed diagram of the Minion system architecture.